A LASER ABLATION SLIT-JET CLUSTER SOURCE FOR HIGH RESOLUTION INFRARED STUDIES

\underline{R. S. WALTERS}, and D. J. NESBITT, JILA, Department of Chemistry and Biochemistry, University of Colorado, Boulder, CO. 80309.

A novel experiment combining a laser ablation cluster source with a slit-jet expansion for the study of metal-molecular radicals by high resolution infrared absorption spectroscopy is presented. This experiment takes advantage of the nearly unlimited capabilities of a laser ablation technique and the high densities, longer path lengths and sub-Doppler characteristics of a slit-jet supersonic expansion. The details of the apparatus including the ablation cluster source and a new high resolution infrared laser spectrometer (2500 cm^{−1} - 5000 cm^{−1}) are discussed.